### v1

# Shot-by-Shot: Film-Grammar-Aware Training-Free Audio Description Generation

Abstract
Our objective is the automatic generation of Audio Descriptions (ADs) for edited video material, such as movies and TV series. To achieve this, we propose a two-stage framework that leverages “shots” as the fundamental units of video understanding. This includes extending temporal context to neighbouring shots and incorporating film grammar devices, such as shot scales and thread structures, to guide AD generation. Our method is compatible with both open-source and proprietary Visual-Language Models (VLMs), integrating expert knowledge from add-on modules without requiring additional training of the VLMs. We achieve state-of-the-art performance among all prior training-free approaches and even surpass fine-tuned methods on several benchmarks. To evaluate the quality of predicted ADs, we introduce a new evaluation measure – an action score – specifically targeted to assessing this important aspect of AD. Additionally, we propose a novel evaluation protocol that treats automatic frameworks as AD generation assistants and asks them to generate multiple candidate ADs for selection.
1 Introduction
In movies and TV series, Audio Descriptions (ADs) are narrations provided for the visually impaired, conveying visual information to complement the original soundtrack. Their purpose is to ensure a continuous and coherent narrative flow, enabling audiences to follow the plot effectively. Unlike video captions, ADs are constrained by length, prioritising the most visually salient and story-centric information, such as character dynamics and significant objects, whilst omitting redundant details like background figures or unchanging locations. Additionally, ADs are typically produced by professional narrators in a specific style and format, ensuring coherence while not interfering with the original audio.
With the advent of Visual-Language Models (VLMs), there has been a growth of interest in automatically generating ADs for both movies [94, 98, 27, 28, 29, 80, 51, 20] and TV material [90, 23]. However, as anyone who has ever watched a movie or read a book about film editing knows – the fundamental unit of edited video material is the shot, not the frame [56, 41, 58]. Shots are used to structure the video material, defining the granularity and temporal context through choices in its scale (close-up, long shot, etc.) and its movement (panning, tracking, etc.). The “film grammar” is then used to convey meaning through specific editing choices on shot transitions (cuts, fades, etc.), durations, and composition (threads, montage, etc.).
Current approaches to automating AD generation, and video large language models (VideoLLMs) in general, are not shot-aware, hindering interpretation of edited video material that has frequent shot transitions [76]. In this paper, our principal objective is to incorporate shot information and editing structure into the AD generation process. To this end, we consider temporal context in terms of shots, and take account of the two key properties of thread structure and shot scale.
Thread structure identifies the sequence of shots captured with the same camera. An example of a scene with multiple simultaneous threads in space-time is shown in Fig. 3. We develop a robust thread clustering method and use its predictions to guide VLMs in understanding shot-wise relationships.
Shot scale typically implies the type of content of the frame. For instance, close-up shots often highlight characters’ facial expressions or gaze interactions, while long shots tend to depict the overall environment and ambience of the scene, as shown in Fig. 4 (top). We leverage this property by developing an off-the-shelf shot scale classifier and employing a scale-dependent prompting strategy to improve the contextual relevance of VLM-generated descriptions.
In this paper, we develop a training-free AD generation approach. Inspired by AutoAD-Zero [90], as illustrated in Fig. 1, we adopt a two-stage pipeline in which a VideoLLM generates dense text descriptions in the first stage, followed by an LLM that produces the final AD outputs from this text. We improve the effectiveness of training-free approaches by incorporating two key improvements based on: (i) shot-based temporal context; and (ii) the shot scale and thread structure, that are both crucial to cinematic composition and understanding.
As well as the challenge of generating ADs, another challenge is how to evaluate the predicted ADs. Apart from conventional metrics [79, 49, 10, 62], several new AD metrics [98, 29, 27] have been proposed, including CRITIC [29] measuring who is mentioned in the AD. However, these metrics fail to emphasise character actions – one of the most critical aspects of ADs. To address this, we introduce an action score, assessing whether the predictions accurately capture the actions described in GT AD, independent of character information.
Moreover, as ADs are time-limited, there is always a choice on which visual aspect (characters, actions, objects, etc.) should be included. Consequently, a single video clip can have multiple equally valid ADs, each highlighting a different aspect. This observation is supported by the inter-rater agreement experiments in AutoAD-III [29] and previous user studies [86]. Therefore, beyond evaluating single AD performance, we assess our framework’s capability as an assistant to generate multiple AD candidates, and report the performance of the selected one AD.
In summary, we propose an enhanced training-free AD generation framework, with the following contributions:
We incorporate shot-based temporal context into AD generation via training-free prompting techniques including shot number referral and dynamic frame sampling.
We develop state-of-the-art methods for thread structure and shot scale predictions, and demonstrate that incorporating predicted film grammar knowledge enhances AD generation.
We improve the current AD evaluation by introducing the character-free action score, and a new assistance-oriented evaluation protocol.
Our approach achieves state-of-the-art performance in training-free AD generation, and furthermore, surpasses fine-tuned models on multiple benchmarks. This is the first time a training-free approach has achieved superior performance to fine-tuned methods.
2 Related work
Audio Description generation. Efforts have been made to curate Audio Description (AD) datasets for both movies [75, 29] and TV series [90], with human annotations sourced from platforms such as AudioVault [2].
For automatic AD generation, prior works [27, 28, 29, 51, 80, 23] fine-tune pre-trained VLMs on AD annotations to produce descriptions in an end-to-end manner. However, these methods face challenges due to limited high-quality AD annotations and the high computational cost of fine-tuning each new backbone. Alternatively, training-free approaches [98, 94, 20, 90] have gained traction for their scalability and flexibility, allowing customised AD output based on official guidelines [1] or specific needs. Yet, these methods still lag in performance, while our approach is the first to achieve results on par with fine-tuned methods.
Instead of limiting the video input to each AD interval, UniAD [80] and DistinctAD [23] fine-tune VLMs with multiple AD clip inputs to incorporate broader temporal context. In contrast, our method systematically extends AD clip to adjacent shots and introduces a training-free approach that enables pre-trained VideoLLMs to better capture localised temporal context.
Film grammar analysis. Prior research has sought to understand film grammar from two major perspectives: (i) intra-shot properties, (ii) shot-wise relationships.
For individual shots, several datasets [69, 6, 35, 40, 73] categorise their characteristics based on camera setups, including shot scales (examples shown in Fig. 4) and camera movements. Correspondingly, various models [69, 17, 52, 46, 78] have been developed for shot type classification.
Regarding shot-wise relationships, a few datasets [6, 12, 63] have been proposed to explore transitions (i.e. cuts) between shots, which have also been leveraged in video content generation [70, 101, 24, 64]. Beyond pairwise shot transitions, research has also investigated longer temporal contexts with thread-based editing structures. Notably, Hoai et al. introduced the Thread-Safe [31] dataset, demonstrating that thread information can enhance action recognition. These structures have also been utilised in video-based face and people clustering [77, 11].
While prior work on film grammar has mainly focused on classification and generation tasks, we specifically utilise shot scale and thread structure information to enhance AD generation in movies and TV series.
Dense video captioning. Dense video captioning is closely related to AD generation. Early works [43, 36, 37, 82, 84] in video captioning typically treat event localisation and captioning as independent stages, whereas more recent approaches [93, 65, 88, 38, 42, 106, 105, 85, 14, 22, 57, 68] integrate these tasks in an end-to-end manner. Video captioning benchmarks cover a range of domains, including cooking [104], actions [43], movies [72], TV series [45], and open-domain settings [74, 54, 102, 96, 92, 30, 16, 13, 91].
Temporal grounding in VLM. To equip conventional VLMs with temporal grounding capability, some studies [33, 87, 48] generate additional data with enhanced temporal information for fine-tuning. A more common approach explicitly incorporates temporal information into inputs, either by inserting temporal tokens [34, 32, 25, 26, 66, 81, 15, 21] or embedding time information within visual tokens [33, 87, 48].
Training-free approaches [67, 103, 89] have also been explored for temporally grounded understanding. Notably, a recent work [89] achieves temporal grounding by overlaying frame numbers as visual prompts. Instead of using uniformly sampled timestamps, we leverage the natural shot structures in movies and TV series, adopting them as fundamental units for training-free temporal referral.
Evaluation for Audio Description. Early AD evaluation often adopts captioning metrics, including CIDEr [79], ROUGE [49], BLEU [62], and METEOR [10], which measure word/n-gram overlap with different weighting strategies. Other metrics focus on semantics, leveraging rule-based scene graph matching (SPICE [5]), word embedding similarity (BERTScore [100], Recall@k/N [27]), or LLM-based evaluation [29, 98].
However, the fixed and distinct set of character names in each video can bias conventional captioning metrics. For instance, TF-IDF [71] weighting in CIDEr assigns high importance to character names. To provide a more independent measure of character names and other AD content, CRITIC [29] assesses character recognition in predicted ADs, while we develop an action-centric metric with minimised dependence on character names.
Most prior works evaluate single AD outputs; however, training-free methods are compatible with generating multiple candidates for selection. While early studies [86] explored this through user studies, we propose a quantitative evaluation protocol to further assess AD generation methods as assistive tools.
3 Training-free AD generation framework
Given a video clip , the task of audio description is to generate a concise narration describing what happens around a given AD interval . In this work, we propose a two-stage framework that leverages VideoLLMs and LLMs to predict ADs in a training-free manner.
In Stage I, we employ a VideoLLM to generate a dense description for the clip based on instructions :
In Stage II, we then prompt an LLM (with instructions ) to extract key information from the dense Stage I description and format it into an AD-style narration :
In this section, we focus on enhancing the visual understanding of the Stage I VideoLLM for edited video material, and make three innovations: In Sec. 3.1, we incorporate shot-based temporal context into Stage I visual inputs. In Sec. 3.2, we leverage the thread structure to enrich cross-shot understanding. Finally, in Sec. 3.3, we incorporate shot-scale awareness into Stage I prompt formuation.
3.1 Leveraging shot-based temporal context
Regarding the visual inputs to VideoLLM, prior works [51, 29] often sample frames that directly correspond to the AD interval . However, this approach can be problematic due to (i) misalignment between the AD interval and the actual timestamps when the action occurs, and (ii) the lack of contextual information from adjacent shots. Therefore, we investigate how incorporating temporal context can enhance the understanding of the video clip.
Structuring temporal context with shots. To obtain more visual context information, instead of simply extending the AD interval by fixed timestamps, we explore a more structured approach that treats shots as the fundamental units.
Specifically, we apply an off-the-shelf shot segmentation model to partition the entire video clip into individual shots. For each AD interval, we first identify the shots that (partially) overlap with it, referred to as “current shots”, as illustrated in Fig. 2. We then consider at most two “past shots” and two “future shots” adjacent to current shots as temporal context. For all shots included, we label them sequentially from past to future with a number starting from “Shot 0”.
Emphasising the targeted AD interval. To ensure that the VideoLLM focuses on describing visual content within the AD interval, we propose two strategies: dynamic frame sampling and shot number referral.
Dynamic frame sampling. As shown in Fig. 2, to emphasise the frames of interest, we adopt denser sampling within the AD interval (red region) and sparser sampling for the surrounding context frames (yellow region). In practice, we specify fixed numbers of frames to be sampled within and outside the AD interval and apply uniform sampling according to these constraints.
Shot number referral. To further enhance the attention towards the current shot content, we label each sampled frame with its shot number (e.g. “Shot 0”) at the top-left. During the formulation of the text prompt, instead of prompting the VideoLLM to “describe what happened in the video clip”, we ask it to “describe what happened in [Shot 2, Shot 3]” (i.e. current shots). Through this visual-textual prompting strategy, we found that the VideoLLM could successfully interpret the meaning of shot numbers and refer to the correct shots.
3.2 Leveraging thread structure
Movies are generally edited such that viewpoints from two or more cameras are intertwined in shot threads, as illustrated in Fig. 3 (top). These interleaved arrangements of shot threads often imply relationships between objects and characters (e.g. gaze interactions) and their 3D arrangement. To leverage this information for Stage I description, we first determine the thread structure using a separate module, then incorporate it into the VideoLLM through prompt guidance.
Thread structure prediction. We develop a training-free method to predict thread clustering. The problem here is to determine if two shots correspond to the same viewpoint or not. Specifically, given two shots (Shot and a later Shot ), we compare the last frame of Shot () with the first frame of Shot (), extracting their DINO features as and .
Inspired by [39], we assess frame-wise dense feature correlations by computing a cost volume between their feature maps:
where and are the normalised -th and -th spatial elements in the respective feature maps. The binary attention mask is set to only if the spatial position of the -th element is within an neighbourhood of the -th element.
We then apply a softmax operation along the last dimension () and find the maximum similarity for each , followed by averaging over all -th elements to obtain a matching score between Shot and Shot :
where denotes the number of feature patches, and is the softmax temperature.
Intuitively, as shown in Fig. 3, this process effectively checks whether each patch () in one shot frame matches with a patch () in the other shot frame at a roughly similar spatial position (i.e., within an neighbourhood).
Finally, we construct an adjacency matrix based on the predicted scores for all possible pairs of Shot and Shot . By thresholding the adjacency matrix (threshold ) and identifying the largest connected components, we cluster the set of shots into multiple threads.
Thread structure injection. Once the thread information is obtained, we inject it into the text prompt for the Stage I VideoLLM. In practice, we conduct this information injection only to clips that exhibit thread structures (i.e. ). For each given thread , we formulate the prompt as: “ share the same camera setup”. This statement implies that the cameras in these shots maintain consistent angles and scales.
Rather than simply providing this information, we further engage the VideoLLM by asking it to explain why the given thread structure is correct. This effectively corresponds to a Chain-of-Thought (CoT) process, enhancing its understanding of these repetitive thread structures.
3.3 Leveraging shot scale information
In movies and TV series, shot scales are often carefully designed during filming or post-editing to implicitly convey information to the audience. Our objective is to use the shot scale to choose what should be included in the Stage I prompts.
Shot scale classification. We first build a shot scale classification network by fine-tuning a pre-trained DINO model. For all current shots, we classify their shot scales into one of five classes, represented by values –, as shown in Fig. 4 (top). We then compute their average to obtain the effective shot scale .
Stage I factor composition. We then leverage the predicted shot scale to determine the factors to include in Stage I instructions. We first consider three fixed factors that form the basis of ADs, namely characters, actions, and interactions. The additional factor can be determined through applying a set of thresholds to , as detailed in Fig. 4 (bottom). For example, for close-up shots (), we would ask the VideoLLM to additionally describe the “facial expression”, whereas for long shots (), a description on the “environment” will be included.
4 Action score
In this section, we introduce a new metric, termed “action score”, which focuses on whether a specific ground truth (GT) action is captured within a predicted Stage I description paragraph or a Stage II AD output. For instance, for the GT action “He dials the phone”, we want the metric to measure the action of “dial the phone”, but not be sensitive to character names and other predicted content. Therefore, this metric is designed to possess two key properties: (i) it is character-free, meaning that the presence of character names has minimal impact on the evaluation, and (ii) it is recall-oriented, without penalising additional action information in the prediction.
Preprocessing of GT actions. For each GT AD, we extract the character-free GT actions by (i) replacing character names with pronouns, and (ii) splitting the AD into subsentences, each containing one action (verb). For example, given a GT AD “Chandler dials the phone, then hurriedly hangs up.”, the extracted GT actions are “He dials the phone.” and “He hurriedly hangs up.”
Hierarchical parsing of predictions. To process the predicted paragraph during evaluation, we first decompose it into individual sentences (light blue), as shown in the left column of Fig. 5. For each predicted sentence, we then perform rule-based dependency parsing to obtain action phrases (deep blue). Then, all extracted sentences and action phrases are collected to form a prediction set. Additionally, for each action phrase, we extract the corresponding verb lemma (i.e. the root form of verbs).
Action score computation. The action score features a combination of semantic-based and verb-matching-based components, as illustrated in the middle and right column of Fig. 5, respectively.
Semantic-based evaluation. For each GT action sentence, we assess its semantic similarity with the set of predicted sentences and phrases. Specifically, we employ a general text embedding (GTE) model to compute sentence-level embeddings for the GT action sentence and each element in the predicted set . By computing cosine similarities and taking the maximum value, the similarity score is defined as .
Verb-matching-based evaluation. In addition to semantic matching, we credit predictions that contain the same verbs as those in GT actions. Practically, we compare the predicted verb lemmas with GT verb lemma and compute binary matching scores . These matching scores are further weighted by the corresponding similarity scores through element-wise multiplication. Finally, we take the maximum of the regulated scores to compute the verb score, i.e. .
To obtain the final action score, we combine the scores from both sources using a weighted average , with and denoting the weighted factors.
5 Experiments
This section begins with datasets and metrics for AD generation in Sec. 5.1, followed by implementation details in Sec. 5.2. Results on film grammar predictions are presented in Secs. 5.3 and 5.4, while Sec. 5.5 analyses human alignment with action scores. AD generation results are detailed in Secs. 5.6, 5.7 and 5.8.
5.1 AD generation datasets and metrics
We evaluate our framework on AD generation datasets for both movies (CMD-AD [29], MAD-Eval [27]) and TV series (TV-AD [90]). In more detail, CMD-AD is constructed by aligning ground truth ADs with the Condensed Movie Dataset (CMD) [9], comprising k ADs (k for training and k for testing) from k movies. MAD-Eval consists of k ADs sampled from movies within LSMDC [72]. On the other hand, TV-AD features k AD annotations from TV series, with its test set sourced from TVQA [44], consisting of k ADs.
5.2 Implementation details
In this section, we provide key implementation details for AD generation and action score computation. For additional information on film grammar predictions and other details, please refer to Appendix A.
Shot detection. To segment the video clip into shots, we use PySceneDetect [3] with the “Adaptive Detection” method, which compares the ratio of pixel changes with the neighbouring frames. On average, each shot spans s, s, s in CMD-AD, TV-AD, and MAD-Eval, respectively.
AD generation setting. During dynamic frame sampling, we select a total of frames, with frames uniformly sampled within and outside the AD interval, respectively. For character recognition, we adopt the same visual-textual prompting method proposed by [90], which applies coloured circles around faces for visual character indication. For simplicity, the visualisations in this paper do not display these circle labels.
Regarding the base models, we use Qwen2-VL-7B [83] as the VideoLLM in Stage I and LLaMA3-8B [53] as the LLM in Stage II. This setup is used as the default unless stated otherwise. Additionally, we explore the framework with the proprietary GPT-4o [60] model for both stages.
Action score evaluation setting. For action score computation, we set the weight factors as = and = . When aggregating the action score results, we first average over multiple GT actions within each GT AD, and then perform global averaging across all AD samples. Moreover, in practice, we find that most action scores are clustered within the range of . To improve clarity, we apply further rescaling as post-processing. For evaluations in this paper, unless otherwise specified, we use the action score to assess the Stage II AD outputs.
| Feature | Frame setup | Precision | Recall | AP | WCP |
| CLIP-L14 CLS | Side | ||||
| DINOv2-L14 CLS | Side | ||||
| DINOv2-g14 CLS | Side | ||||
| DINOv2-g14 spatial | Mid | ||||
| DINOv2-g14 spatial | All | ||||
| DINOv2-g14 spatial | Side |
5.3 Thread structure prediction
We evaluate thread structure prediction on Thread-Safe [31], which consists of approximately k video clips collected from TV series. Each video clip contains a multi-shot scene with corresponding thread clusters manually annotated.
For evaluation, we first construct an adjacency matrix from the GT clusters and extract binary GT labels from the off-diagonal entries, where each label indicates whether a given pair of shots belong to the same cluster. We then compute the Average Precision (AP) between the predicted relationships and the ground truth , as well as report the precision and recall values. Additionally, we directly compare the predicted clusters with ground truth clusters by reporting the weighted clustering purity (WCP) [77].
Tab. 1 verifies the choice of DINOv2-g14 as the feature extractor for frame pair comparison. Compared to abstract CLS tokens, dense spatial matching achieves higher AP in frame pair relationship prediction and higher WCP in thread clustering. Additionally, we observe that using the temporally closest frames from two shots (“Side”) leads to a noticeable performance improvement. This can be attributed to the continuous story flow across shots within the same thread.
5.4 Shot scale classification
Following prior work on shot scale classification [52, 46], we use the MovieShots [69] dataset, which consists of k shots (train:val:test = ::) collected from over k movie trailers. We follow its definition of shot scales, categorising shots into five classes ranging from extreme close-up to long shots, as illustrated in Fig. 4 (top). To evaluate the model performance, we report classification accuracy and Macro-F1 [46] scores on the MovieShots test set.
Since previous state-of-the-art methods on shot scale classification are not open-sourced, we develop a new network by fine-tuning DINOv2, achieving superior performance over prior approaches that rely on additional optical flow or SAM-based mask inputs, as demonstrated in Tab. 2.
| Metric | Input | Accuracy | Macro-F1 |
| ViViT [7] | RGB | ||
| SGNet [69] | RGB + Flow | ||
| Lu et al. [52] | RGB + mask | ||
| Li et al. [46] | RGB | ||
| Ours | RGB |
| Metric | Paragraph | Sentence | ||
| Pearson | Spearman | Pearson | Spearman | |
| CIDEr [79] | ||||
| ROUGE-L [49] | ||||
| METEOR [10] | ||||
| BLEU-1 [62] | ||||
| SPICE [5] | ||||
| BERTScore [100] | ||||
| LLM-based (LLaMA3.1-70B [53]) | ||||
| LLM-based (GPT-4o [60]) | ||||
| Action Score (w/o verb matching) | ||||
| Action Score (w verb matching) | ||||
5.5 Human alignment with action scores
Action scores aim to evaluate whether a GT action is captured within a predicted description, making them recall-oriented. Such descriptions can be in the form of paragraphs (Stage I descriptions) or single sentences (Stage II ADs). To assess whether action scores align with human judgments, we create a dataset containing pairs of predicted descriptions and GT actions. For each GT action, human annotators manually annotate the quality of predictions into based on the relevance towards GT action ranging from “unrelated” (0) to “exact matching” (3).
Comparison with other metrics. Next, we use the human-annotated scores as a reference to compare different metrics in terms of human agreement (measured by correlations), as reported in Tab. 3. Additionally, we consider an LLM-based metric to predict scores, following the same scoring criteria as human annotations. In general, the action score achieves the best correlations with human annotations. Note, it is also more efficient than LLM-based metrics, with s compared to s per prediction evaluation. For more details regarding this human agreement study, please refer to Appendix D.
| Exp. | Temporal context | Frame sampling | Shot label | CMD-AD | TV-AD | ||||
| CIDEr | CRITIC | Action | CIDEr | CRITIC | Action | ||||
| A | |||||||||
| B | shot | Dyn. | Shot num. | ||||||
| C | shots | Dyn. | |||||||
| D | shots | Dyn. | Frame box | ||||||
| E | shots | Uni. | Shot num. | ||||||
| F | shots | Dyn. | Shot num. | ||||||
| Thread structure | CMD-AD subset | TV-AD subset | ||||
| CIDEr | CRITIC | Action | CIDEr | CRITIC | Action | |
| Stage I factors | CMD-AD | TV-AD | ||||
| CIDEr | CRITIC | Action | CIDEr | CRITIC | Action | |
| Base | ||||||
| Base + Face (AutoAD-Zero) | ||||||
| Base + Obj. | ||||||
| Base + Env. | ||||||
| Base + Face + Obj. + Env. | ||||||
| Scale-dependent (Ours) | ||||||
| Method | VLM | LLM | Training -free | Propriet. model | CMD-AD | TV-AD | ||||||||
| CIDEr | CRITIC | Action | R@1/5 | LLM-AD-Eval | CIDEr | CRITIC | Action | R@1/5 | LLM-AD-Eval | |||||
| AutoAD-II [28] | CLIP-B32 | GPT-2 | ✗ | ✗ | ||||||||||
| AutoAD-III [29] | EVA-CLIP | LLaMA2-7B | ✗ | ✗ | ||||||||||
| DistinctAD [23] | CLIP-B16 | LLaMA3-8B | ✗ | ✗ | ||||||||||
| Video-LLaMA [99] | Video-LLaMA-7B | ✗ | ||||||||||||
| VideoBLIP [95] | VideoBLIP | ✗ | ||||||||||||
| AutoAD-Zero [90] | VideoLLaMA2-7B | LLaMA3-8B | ✗ | |||||||||||
| AutoAD-Zero [90] | Qwen2-VL-7B | LLaMA3-8B | ✗ | |||||||||||
| Ours | Qwen2-VL-7B | LLaMA3-8B | ✗ | |||||||||||
| AutoAD-Zero [90] | GPT-4o | GPT-4o | ||||||||||||
| Ours | GPT-4o | GPT-4o | ||||||||||||
| Method | VLM | LLM | Training -free | Propriet. model | MAD-Eval | |||||
| CIDEr | R@5/16 | Rouge-L | SPICE | METEOR | BLEU-1 | |||||
| ClipCap [55] | CLIP-B32 | GPT-2 | ✗ | ✗ | ||||||
| CapDec [59] | ✗ | ✗ | ||||||||
| AutoAD-I [27] | CLIP-B32 | GPT-2 | ✗ | ✗ | ||||||
| AutoAD-II [28] | CLIP-B32 | GPT-2 | ✗ | ✗ | ||||||
| AutoAD-III [29] | EVA-CLIP | LLaMA2-7B | ✗ | ✗ | ||||||
| MovieSeq [51] | CLIP-B16 | LLaMA2-7B | ✗ | ✗ | ||||||
| DistinctAD [23] | CLIP-B16 | LLaMA3-8B | ✗ | ✗ | ||||||
| UniAD [80] | CLIP-L14 | LLaMA-8B | ✗ | ✗ | ||||||
| Video-LLaMA [99] | Video-LLaMA-7B | ✗ | ||||||||
| Video-BLIP [95] | Video-BLIP | ✗ | ||||||||
| AutoAD-Zero [90] | VideoLLaMA2-7B | LLaMA3-8B | ✗ | |||||||
| AutoAD-Zero [90] | Qwen2-VL-7B | LLaMA3-8B | ✗ | |||||||
| Ours | Qwen2-VL-7B | LLaMA3-8B | ✗ | |||||||
| VLog [4] | BLIP-2 + GRIT | GPT-4 | ||||||||
| MM-Vid [50] | GPT-4V | |||||||||
| MM-Narrator [98] | Azure API + CLIP-L14 | GPT-4V | ||||||||
| MM-Narrator [98] | Azure API + CLIP-L14 | GPT-4 | ||||||||
| LLM-AD [20] | GPT-4V | |||||||||
| AutoAD-Zero [90] | GPT-4o | GPT-4o | ||||||||
| Ours | GPT-4o | GPT-4o | ||||||||
5.6 AD generation – evaluation of components
Shot-based temporal context. We investigate different setups for leveraging temporal context in AD generation, as shown in Tab. 4, leading to the following observations: (i) Expanding the temporal context range noticeably boosts the performance, with gains saturating around “ shots” (Exp. A, B, and F); (ii) “Shot number referral” is the most effective strategy for outlining the current shot. (Exp. C, D, and F); (iii) Dynamically sampling the current shots at a higher frame rate boosts AD generation (Exp. E and F). The latter two improvements can be attributed to a more efficient focus on the visual content around the targeted AD interval.
Thread structure injection. After extending the context information with neighbouring shots, we further enhance the VideoLLM’s understanding by incorporating thread structures. Note that this guidance is applied only to video sequences exhibiting thread structures. As shown in Tab. 5, this additional information improves AD generation performance across both datasets.
Scale-dependent Stage I factors. Tab. 6 explores the impact of different Stage I factors on final AD performance. Using shot scales as guidance for Stage I factor formulation (i.e. scale-dependent) not only outperforms configurations with single fixed factors but also surpasses the case where all factors are included in Stage I. This could be attributed to that the scale-dependent description contains more relevant and less redundant information, enabling more efficient AD extraction in Stage II.
5.7 AD generation – comparison with SotA
Tab. 7 provides a comprehensive summary of AD generation performance on CMD-AD and TV-AD, comparing across training-free methods with and without proprietary models, as well as models fine-tuned on human-annotated ADs. Notably, with the same base model setup (Qwen2-VL-7B + LLaMA3-8B), our training-free framework significantly outperforms AutoAD-Zero, primarily due to the usage of temporal context and film grammar information. By incorporating the more powerful GPT-4o models, our performance scales up further, surpassing even existing fine-tuned models.
Regarding the MAD-Eval results, as shown in Tab. 8, our framework achieves state-of-the-art performance among all training-free approaches, also demonstrating competitiveness with fine-tuned models.
Qualitative visualisations. Fig. 6 presents several qualitative examples, where the top two cases illustrate how temporal context information aids in identifying key objects.
In the bottom-left example, prior methods fail to associate characters, leading to the omission of the man (Alonzo). In contrast, our method recognises the thread structure (i.e. [Shot 0, Shot 2], [Shot 1, Shot 3]), which guides the correct prediction of the man’s gaze direction towards the lying woman.
The bottom-right example highlights the effectiveness of scale-dependent Stage I factor formulation. AutoAD-Zero, designed to query characters, actions, interactions, and facial expressions, sometimes overlooks environmental details. Our method, in contrast, correctly identifies the shot as a long shot and instructs the VideoLLM to incorporate environmental context, resulting in more accurate scene descriptions. For more visualisations, please refer to Appendix E.
| Method | Candidate sampling | CMD-AD | TV-AD | ||||
| CIDEr | CRITIC | Action | CIDEr | CRITIC | Action | ||
| Ours | Single AD (Ref.) | ||||||
| Ours | Indep. output ( = = ) | ||||||
| Ours | Indep. output ( = = ) | ||||||
| AutoAD-Zero [90] | Joint output | ||||||
| Ours | Joint output | ||||||
5.8 Assisted AD generation
The subjective nature of AD sets a practical limit on metric scores, lower than the theoretical maximum, because human annotators often provide different but valid descriptions. Therefore, beyond enforcing generating a single AD sentence, we also consider employing our framework as an assistant to produce multiple candidate AD sentences.
To standardise such a protocol, we consider five candidate ADs generated by an assistant and employ an expert to select the best one. To effectively benchmark performance against existing GT ADs, we define the “expert” as an automatic selection mechanism that chooses the candidate with the highest average CIDEr and action score.
To develop an AD generation assistant, we fix the Stage I dense descriptions and explore generating multiple candidates in Stage II. This can be achieved by either running Stage II independently five times (termed the “independent output” setup) or generating five AD outputs simultaneously within a single run (termed the “joint output” setup). As observed in Tab. 9, the assistant-based setup significantly improves upon the single AD performance, highlighting the potential of training-free methods in effectively capturing the desired content for AD generation. Within the “independent output” setup, increasing the randomness of sampling (i.e. higher and ) enhances the quality of the selected AD, owing to greater candidate diversity. Meanwhile, the “joint output” setup achieves superior performance, which could be attributed to reduced information redundancy across the simultaneously generated ADs. For additional visualisations, discussions, and detailed text prompts, please refer to Appendix E.
6 Discussion – summary and limitations
We have demonstrated the benefit of shot-based context and film grammar awareness in AD generation – our training-free two-stage framework achieves state-of-the-art performance among all training-free counterparts, even surpassing fine-tuned models on multiple benchmarks.
The current framework has two main limitations: (i) the performance depends on the base VideoLLM, which may occasionally hallucinate details inconsistent with the visual content; and (ii) story-level context is not incorporated into the AD generation process. These limitations could potentially be addressed in future work by improving visual grounding and extending the visual and textual context to include the plot.
Acknowledgements.
This research is supported by the UK EPSRC Programme Grant Visual AI (EP/T028572/1), a Royal Society Research Professorship RPR1191132, a Clarendon Scholarship, and ANR-21-CE23-0003-01 CorVis.
Appendix
This appendix is organised as follows:
Implementation details: In Appendix A, we provide additional details on film grammar prediction, action score evaluation settings, and the exact text instructions.
Evaluation metrics for AD generation: In Appendix B, we elaborate on the key metrics used to measure AD quality.
Additional experimental results: In Appendix C, we provide additional results for AD generation.
Human alignment with action scores: In Appendix D, we provide a thorough description of the human agreement study for action scores, including curated test sets, a correlation analysis, and an inter-rater agreement study.
Qualitative visualisations: Appendix E includes more detailed visualisations, along with an in-depth analysis of failure cases and assisted AD generation.
Appendix A Implementation details
Thread structure prediction setting. To predict thread structure in a zero-shot manner, we resize video frames to p and employ DINOv2 [61] ViT-g/14 w. reg to extract spatial features. During matching score prediction, we consider a mask neighbourhood for cost volume computation and set the softmax temperature to . To determine the relationship between each pair of shots, we apply a threshold of to the matching score .
Shot scale classification setting. For shot scale classification, we fine-tune the last layers of the DINOv2 [61] ViT-B/14 on the MovieShots training set. During evaluation, we use the middle frame of each shot as input.
The averaged shot scale of the current shots (i.e. the effective shot scale ) is used to guide the incorporation of additional factors in Stage I prompts. Specifically,
Action score evaluation setting. To obtain character-free GT action sentences, we employ LLaMA3.1-70B [53] for pre-processing in two steps: (i) Character information removal: Character names are replaced with appropriate pronouns using the LLM with the prompt provided in Algorithm 3; (ii) Action sentence extraction: Each AD sentence is split into subsentences, each containing a single action. To achieve this, the LLM is prompted with instructions in Algorithm 4.
During hierarchical prediction parsing, we use spaCy111https://spacy.io/models/en to extract action phrases and corresponding verb lemmas from predicted sentences.
During similarity score computation, to extract sentence embeddings, we apply gte-Qwen2-7B-instruct [47], which supports optional text prompt input as guidance. We set the prompt to: “Retrieve relevant passages that involve similar actions, with particular focus on the verbs.”, further emphasising actions and verbs during similarity matching.
Additionally, to establish the LLM-based baseline, we define evaluation criteria as outlined in Algorithm 5 and use them to prompt LLaMA-3.1-70B and GPT-4o.
GPT-4o setup for AD generation. For both stages, we use gpt-4o-2024-08-06 [60] as the base model. For visual token extraction, the “detail” parameter is set to “low”.
Text instructions for AD generation. The prompts for AD generation are provided in Algorithms 1 and 2 for Stage I and Stage II, respectively. The Stage I prompt is designed for both Qwen2-VL and GPT-4o, while the Stage II prompt is tailored for LLaMA3 and GPT-4o.
| Stage I VideoLLM | Stage II LLM | CMD-AD | TV-AD | ||||||
| CIDEr | CRITIC | Action (Stage I) | Action (Stage II) | CIDEr | CRITIC | Action (Stage I) | Action (Stage II) | ||
| Qwen2.5-VL-7B [8] | LLaMA3-8B [53] | ||||||||
| VideoLLaMA3-7B [97] | LLaMA3-8B [53] | ||||||||
| InternVL2.5-8B [18] | LLaMA3-8B [53] | ||||||||
| Qwen2-VL-7B [83] | LLaMA3-8B [53] | ||||||||
[…]
Appendix B Evaluation metrics for AD generation
CIDEr [79] measures text similarity by computing a weighted word-matching score, emphasising n-gram overlap while accounting for term frequency and importance through TF-IDF [71] weighting.
Recall@k/N [27] is a retrieval-based metric that evaluates whether predicted texts can be distinguished from their temporal neighbours. Specifically, for each predicted AD, it checks whether the AD can be retrieved at a top-k position within a neighbourhood of N ADs. Following prior work [29, 23], we report Recall@1/5 on CMD-AD and TV-AD, and Recall@5/16 on MAD-Eval.
LLM-AD-Eval [29] employs LLM agents (LLaMA3-8B [53] LLaMA2-7B [19]) as evaluators to compare ground truth ADs with predictions, generating a matching score ranging from 1 (lowest) to 5 (highest).
CRITIC [29] measures the accuracy of character names in predicted ADs. It first resolves character ambiguity in GT ADs by applying a coreference model to replace pronouns with corresponding character names. During evaluation, the intersection-over-union (IoU) of predicted and ground truth character names is computed.
Action Score, as detailed in Sec. 4, evaluates the quality of predicted actions (i.e. verbs, object nouns, etc.) while minimising the influence of character name variations. Throughout this paper, unless otherwise specified, we use the action score to assess Stage II AD outputs, i.e. “Action” refers to “Action (Stage II)”.
For MAD-Eval, we additionally report the performance on conventional metrics including ROUGE-L [49], SPICE [5], METEOR [10], and BLEU-1 [62].
| Thread structure | Stage I VideoLLM | TV-AD subset | ||
| CIDEr | CRITIC | Action | ||
| ✗ | Qwen2.5-VL-7B [8] | |||
| Qwen2.5-VL-7B [8] |
1.2 |
0.6 |
0.0 |
[…]
| VideoLLaMA3-7B [97] |
0.6 |
1.0 |
0.8 |
[…]
| InternVL2.5-8B [18] |
1.2 |
1.6 |
1.3 |
[…]
| Qwen2-VL-7B [83] |
1.9 |
0.7 |
0.3 |
Appendix C Additional experimental results
Different VideoLLMs for Stage I. Tab. A1 compares our AD generation performance using different open-source VideoLLMs in Stage I, validating our choice of Qwen2-VL-7B as the default model.
Beyond the Stage II action scores presented in the main text, we also report Stage I action scores as a direct indicator of dense description performance. In general, Stage I action scores are noticeably higher than their Stage II counterparts, suggesting that some ground truth actions are captured in dense descriptions but are not selected for the final AD outputs. This further supports the validity of our assisted AD generation protocol, where multiple candidate ADs with different actions are extracted from dense descriptions and await further selection.
Thread structure injection for different VideoLLMs. We investigate how injecting thread structure into different open-source VideoLLMs affects AD generation. Specifically, we evaluate performance on TV-AD, which contains a large proportion of thread-structured video clips. As shown in Tab. A2, incorporating thread information leads to general performance boosts across various VideoLLMs.
Repeated AD generation. Tab. A3 presents the results of repeated AD generation experiments, demonstrating consistency across different runs. In particular, the CRITIC results exhibit the highest variance, followed by CIDEr, while the action score remains relatively stable across repeated experiments.
| Exp. | CMD-AD | TV-AD | ||||
| CIDEr | CRITIC | Action | CIDEr | CRITIC | Action | |
| 1 | ||||||
| 2 | ||||||
| 3 | ||||||
| 4 | ||||||
| 5 | ||||||
| Mean | ||||||
| STD | ||||||
| Score | Note | GT action | Prediction |
| 0 |
Unrelated:
Completely irrelevant in actions, objects, environments, etc. |
He indicates that he was just joking. | The video begins with a man sitting under an umbrella in the rain, wearing a dark jacket with red accents and a blue shirt underneath. The scene transitions to show him lying on the ground, wet from the rain, with his arms outstretched and his eyes closed. Another person, dressed in a white shirt and black vest, reaches towards him from above. |
| His new companion cuddles back up to him. | The flames start to die down. | ||
| 1 |
Loosely related:
Semantically or causally relevant actions; Similar objects, environments, etc. |
He points the remote at the word’s main menu. | 1. Main characters: Morty and Michael Newman. 2. Actions: Both characters are standing in a room with futuristic digital screens and symbols around them. They appear to be engaged in an activity involving the screens, possibly interacting with them or using devices. 3. Character-character interactions: The two characters seem to be focused on the screens, possibly discussing or collaborating on something related to the content displayed. |
| They step away from the window. | The four of them leave the cabin and go to their horses tethered to a rail at the side. | ||
| 2 |
Similar meaning:
Similar verb, objects, environments, etc. |
He dives into an Olympic-sized swimming pool. | The movie clip shows a young boy running out of a house wearing only a diaper. He then jumps over a fence and runs across a lawn. The boy continues running and jumps into a swimming pool. Throughout the video, the boy’s movements are energetic and playful, and he appears to be having fun. |
| She carries a laundry basket outside the house. | Caroline brings some washing into the garden. | ||
| 3 |
Exactly same match:
Exactly the same verb; Same/similar objects, environments, etc. |
He lights a cigar. | In the movie clip, a man is seen sitting at a table, while another man enters the room and approaches him. The second man takes out a cigarette and lights it, while the first man watches him. The scene is set in a dimly lit room with a table and chairs in the foreground, and a window in the background. |
| She stares glumly at the night sky. | At the palace, Jasmine wanders out into her balcony and stares up at the stars. |
Appendix D Human alignment with action scores
To assess how the proposed action score aligns with human knowledge, we curate a test set that measures the correlation between action scores and human scoring.
Scoring criteria. The human-annotated scores measure whether the ground truth (GT) action is described in the descriptions, ranging from to based on the following scoring criteria:
Score 0 - GT action is unrelated to any action in PD
Score 1 - GT action is loosely related to an action in PD
Score 2 - GT action is similar in meaning to an action in PD
Score 3 - GT action exactly matches with an action in PD, using the same verb
where PD stands for Predicted Description. Additional guidelines and scoring examples are provided in Tab. A4.
Test Set Formulation. We construct two test sets, namely the “paragraph set” and the “sentence set,” corresponding to the scoring of (Stage I) dense descriptions and (Stage II) AD sentences, respectively.
The paragraph set consists of ground truth (GT) ADs, each paired with a predicted paragraph. In total, around character-free GT actions are extracted, with each action-paragraph pair manually annotated by five workers using the – scoring scale, as described in the previous section. For ADs containing multiple GT actions, the final human score is obtained by averaging the manually assigned scores across different actions, meaning the resultant score may not always be an integer.
The sentence set contains GT ADs with approximately actions. Given the video clip described by the GT AD, instead of generating AD predictions from a VLM, we use a human-narrated AD from alternative sources for the same clip as the prediction. Similarly, the final human score is computed by averaging the scores across different actions within each GT AD.
| tIoU | #movies | #AD pairs | CIDEr | R@1/5 | Action | LLM-AD-Eval |
Correlation between human scoring and metrics. We plot the human-annotated scores against the scores reported by each metric, as shown in Fig. A1. Most conventional metrics (blue and pink) fail to align with human evaluations of action predictions. In contrast, both LLM-based metrics and our action scores effectively assess the quality of action predictions in AD sentences (Fig. A1, bottom).
However, when evaluating longer paragraphs against the GT action, LLM-based metrics struggle, whereas action scores maintain a high correlation with human judgments (Fig. A1, top). Quantitative results on human-metric correlations are provided in Tab. 3.
Inter-rater analysis for action score. To obtain an upper bound on action scores for predicted AD sentences, we measure the agreement between two versions of human-annotated ADs for the same movie clip at different temporal IoUs (i.e. inter-rater agreement [29]). As shown in Tab. A5, the action score increases monotonically as the temporal IoU increases.
Appendix E Qualitative visualisations
Fig. A2 presents additional visualisations for CMD-AD and TV-AD, comparing our method against other approaches.
Example failure cases are illustrated in Fig. A3. The top example highlights a hallucination issue in the prediction by Qwen2-VL + LLaMA3 (Ours), where the VideoLLM model misidentifies a “gift” as “candy” and infers an incorrect action of “eating,” which does not occur. When a stronger base model is used (Ours* with GPT-4o), this issue is mitigated.
Additionally, the current method struggles to incorporate broader (story-level) contextual understanding, as demonstrated in the bottom example of Fig. A3. Specifically, the model fails to describe the return of an insect and instead focuses on the sleeping woman. This limitation could potentially be addressed by incorporating more abstract information, which could be extracted from existing movie plots or summarised from a longer temporal context.
Visualisation of intermediate outputs. Fig. A4 provides more detailed visualisations, including intermediate results on thread structure and shot scale predictions, as well as Stage I descriptions. It also illustrates how the predicted shot scales influence the formulation of Stage I factors.
Visualisation for assisted AD generation. Fig. A5 presents additional examples with Stage I descriptions, from which multiple AD candidates are extracted. Among the five AD predictions, the one that best aligns with the ground truth (based on the averaged CIDEr and action score) is highlighed.
In the top example, multiple actions are present in the Stage I dense description (e.g. “kiss”, “hand on neck”, “eyes closed”, etc.), resulting in AD candidates that differ in both subjects and actions. In contrast, the middle example contains fewer actions (e.g. “shoot blue energy” and “look”/“observe”). In this case, the AD candidates primarily vary in style, such as changes in subjects or sentence structures, providing varied options for selection.